Solve the following pair of linear equations by the substitution method 
x + y = 14
 x – y = 4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given,
x + y = 14
and x – y = 4 are the two equations.
From 1st equation,
we get,
x = 14 – y
Now, substitute the value of x in second equation to get,
(14 – y) – y = 4
14 – 2y = 4
2y = 10
y = 5
By the value of y, we can now find the exact value of x;
∵ x = 14 – y
∴ x = 14 - 5
Or x = 9
Hence, x = 9 and y = 5 is the solution.

Thanks